OpenStudy (anonymous):

What is the ratio of x to y? 3/2=2x/2x+y i give medal

OpenStudy (anonymous):

please help :(

OpenStudy (anonymous):

are u familiar with cross multiplication ?

OpenStudy (anonymous):

-.- yes

OpenStudy (anonymous):

so wht do u get ? after cross multiplication

OpenStudy (anonymous):

nuthing cuz it is a lette rproblem duuh

OpenStudy (anonymous):

3*(2x + y) = 2* (2x) ok ? as if we have a/b = c/d ==> a*d = c * b

OpenStudy (anonymous):

so just distribute 3 in bracket

OpenStudy (anonymous):

... u talk like u black... and im not freindly with them... i almost got drowned by one and im 12

OpenStudy (anonymous):

and white

OpenStudy (anonymous):

just try to use this rule of cross multiplication a/b = c/d ==> a*d = c * b and simplify it . to get x/y on left hand side .. hope this will help u :)